No, because it is smaller than the minimum usual value (10) Heights of men on a baseball team have a bell-shaped distribution with a mean of 185 cm and a standard deviation of 5 cm. Using the empirical rule, what is the approximate percentage of the men between the following values? Please show your work! A. 4. % of the men are between 180 cm and 190 cm. B. % of the men are between 170 cm and 200 cm.
Answer #1

Let X:Heights of men on basket ball team

Z=\frac{X-\mu}{\sigma}

A)To find

P(180<X<190)=P(\frac{180-185}{5}<Z<\frac{190-185}{5})=P(-1<Z<1)

By using impirical rule, 68% of data falls within the first standard deviation from the mean.

Therefore,

P(180<X<190)=0.68

B)To find

170 185 200 185 P(170 < X < 200) = P(

By using impirical rule, 99.7% of data falls within three standard deviation from the mean.

Therefore,

170 < X < 200) = 0.997